# Assessing differences between local galaxy dust attenuation and point source extinction within the same environments

### v1

@@ Abstract
Methods. We employ SKIRT, a radiative transfer code, to simulate observations of point sources embedded in dust clouds, as well as spiral and elliptical galaxies. We examine various galaxy morphologies, inclinations, and instrument apertures.
Results. We find that in galaxies the scattering of light into the line of sight and the presence of sources at different depths within the galaxy make attenuation fundamentally different from extinction. For a medium with intrinsic extinction curve slope $R_{V}=3.068$, we recover effective attenuation curve slopes $R_{V_{eff}}$ ranging from $0.5$ to $7$, showing that the two quantities are not analogous, even for local resolved observations. We find that $R_{V_{eff}}$ greatly depends on dust density, galaxy morphology, and inclination, the latter being the most significant. A single simulated galaxy, viewed from different angles, can qualitatively reproduce the well-known relation between attenuation strength $A_{V_{eff}}$ and $R_{V_{eff}}$ observed for star-forming galaxy samples. An increase in dust density leads to higher $R_{V_{eff}}$ across all inclinations, which, assuming a correlation between stellar mass and dust density, explains the increase in $R_{V_{eff}}$ with mass observed in star-forming galaxies. However, we are unable to explain the differences in $R_{V_{eff}}$ between star-forming (higher $R_{V_{eff}}$) and quiescent (lower $R_{V_{eff}}$) high-mass galaxies.
Conclusions. We conclude that highly attenuated regions of simulated face-on galaxies yield $R_{V_{eff}}$ within a $10\%$ error of the intrinsic extinction curve slope of the medium, allowing for the distinction of different dust types. For edge-on spirals, however, the median $R_{V_{eff}}$ for low $A_{V_{eff}}$ regions appears to better approximate the extinction curve slope.
@@ 1 Introduction
Type Ia supernovae (SNe Ia) have long served as fundamental distance indicators for observational cosmology. They have contributed to the measurement of the accelerated expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999) and to the constraint of various cosmological parameters, such as the Hubble constant $H_{0}$ (e.g. Dhawan et al., 2018; Riess et al., 2022; Galbany et al., 2023). However, when comparing the values of $H_{0}$ obtained via SNe Ia and other distance indicators of the distance ladder with those obtained from early-time probes, a tension of $\sim 5\sigma$ has been reported (Planck Collaboration et al., 2020; Riess et al., 2020). Although several cosmological model extensions have been proposed as the solution to this tension (e.g. Karwal & Kamionkowski, 2016; Zhao et al., 2017; Raveri, 2020), the possibility that some systematic error is, at least partially, responsible has not yet been fully discarded. In fact, recent results show that by using an inverse distance ladder where SN distances are calibrated via baryon acoustic oscillations (BAO) data, an $H_{0}$ compatible with early-time measurements can be recovered (Camilleri et al., 2025), pointing to the fact that, from the SN Ia point of view, the tension might originate in the standard local distance ladder (Wojtak & Hjorth, 2024).
@@
The process of SN Ia calibration for cosmology relies on a number of empirical corrections based on observed color-luminosity and light-curve shape-luminosity (Phillips, 1993; Tripp, 1998) relations, which have usually been assumed to be universal (Guy et al., 2010). However, SNe Ia originating in high-mass galaxies have been shown to be more luminous than those originating in low-mass galaxies after the shape and color corrections (Sullivan et al., 2010; Kelly et al., 2010; Lampeitl et al., 2010). This effect is known as the “mass step” and it could hint toward the existence of a systematic bias in the calibration, which could also potentially influence SN Ia cosmological results.
@@
One likely source for the mass-step as a systematic in calibrated SN luminosities is the incomplete treatment of the previously mentioned color-luminosity relation. On top of its assumption of universality, it also combines two different physical effects in the same linear relation: i) intrinsic color-luminosity relations, i.e., fainter SNe Ia are redder; and ii) extrinsic dust effects, i.e., more dust extinction leads to fainter magnitudes and redder SNe through selective absorption. González-Gaitán et al. (2021) propose the existence of two separate populations of SNe Ia as an explanation for the mass-step, with each having different intrinsic color-luminosity properties and dust environment characteristics. This is supported by Grayling et al. (2024), who find evidence for both intrinsic and dust environment differences in SN Ia from different host galaxy mass populations. Brout & Scolnic (2021) suggest that host galaxy dust differences are solely responsible for the effect and define two separate $R_{V}$ populations for low- and high-mass galaxies, finding significantly different values for each one and an overall reduction in the mass-step. Similarly, Wiseman et al. (2022) find that using separate $R_{V}$ populations for younger and older galaxies results in a similar improvement, while Johansson et al. (2021) recover no mass-step when using an individual best-fit $R_{V}$ for each SN Ia in the calibration. Duarte et al. (2023) find that SN host galaxy properties can be used to define a “dust-step”, which nevertheless fails to fully account for the mass-step. It has also been shown that apparent galactic dust properties can vary considerably between different galaxies and that they strongly correlate with environmental properties such as the stellar mass and the stellar age of the population in question (e.g., Garn & Best, 2010; Zahid et al., 2013; Salim et al., 2018; Duarte et al., 2023). As such, assuming a standard universal dust effect for SN Ia might prove inadequate in correcting the color-luminosity relation and could ultimately be, at least partially, responsible for the observed mass-step.
@@
In an effort to improve SN Ia calibration, many methodologies have been recently proposed (e.g. Brout & Scolnic, 2021; Wiseman et al., 2022; Duarte et al., 2023), relying on the assumption that galactic dust properties can be used as an adequate proxy for SN Ia extinction. These approaches, while in some cases presenting a degree of success in reducing the Hubble residuals for the calibrated SN, might ultimately prove to be flawed, given that attenuation and extinction are not the same. While they appear similar, the reddening laws describing the effects of dust on SN and other point source observations, also known as extinction laws, are fundamentally different from the reddening laws affecting integrated light observations such as galaxies, also known as attenuation laws (Krügel, 2009; Salim & Narayanan, 2020). This also means that the parameters usually used to describe dust effects, while sharing the same nomenclature in most of the literature, must be interpreted with care and in different ways.
@@
As light travels through the interstellar medium (ISM) and interacts with dust particles, it experiences absorption and scattering effects, which are especially prevalent for blue wavelengths (Calzetti et al., 1994). These effects lead both to a dimming and a reddening of the original spectral energy distribution, which can be parameterized by an empirical reddening law (e.g. Cardelli et al., 1989; Fitzpatrick, 1999; Calzetti, 2001).
@@
For a point source, the total extinction for the $V$ band, $A_{V}$, which serves as a normalization for the extinction curve, describes the dust column density in the optical path of the observed light. Likewise, the color excess for the $B-V$ color index, $E(B-V)$, refers only to the reddening experienced along that single line of sight, which means that the total-to-selective extinction $R_{V}=A_{V}/E(B-V)$ is only a function of the optical properties of the dust grains themselves (Draine, 2011).
@@
For an extended stellar population or a galaxy, this picture is complicated as a result of two primary factors. On the one hand, the necessity of using a larger aperture for galaxy observations makes the light scattered back into the line of sight a relevant component of the overall observed flux, with the strength of this scattered component being dependent on the morphology and inclination of the galaxy, as demonstrated by Chevallard et al. (2013). Those authors show that galaxy inclination can drastically impact attenuation properties, meaning that identical galaxies viewed from different angles can produce wildly different attenuation curve strengths and slopes. On the other hand, the interspersion of stars and dust lanes leads to varying column densities inside the galaxy, along with different stellar light contributions. Viaene et al. (2017) found that the thicker the dust lane, the more the emission is dominated by stars located in the outermost regions of the galaxy, which are not meaningfully attenuated.
@@
Although this has been long known, it has only recently permeated the SN field. Chevallard et al. (2013) and Viaene et al. (2017) show that a deeper knowledge of the underlying galaxy geometry is needed to properly describe dust attenuation. Duarte et al. (2023) show that the SN Ia calibration returns much higher Hubble residuals when galactic dust properties are used to directly constrain the color luminosity correction. Popovic et al. (2024) conclude that dust in a SN line of sight cannot be accurately described using its host galaxy dust properties.
In this work we present a comprehensive analysis of how point source extinction relates to galactic attenuation. We rely on radiative transfer simulations to systematically study how factors such as instrument aperture and galaxy inclination, morphology and dust composition influence the recovered values of $A_{V_{eff}}$ and $R_{V_{eff}}$ and how these relate to the real optical properties of the medium. More specifically, we seek to understand under which conditions, if any, a host galaxy dust measurement can be said to truly reflect the extinction experienced by a local point source, such as a SN, and therefore be used as its proxy. We use the radiative transfer code SKIRT (Camps & Baes, 2015, 2020) to simulate observations for both point sources and a variety of galaxies (spiral and elliptical), observed from multiple angles and with differing apertures, the details of which are presented in Section 2. In Section 3, we present and discuss our simulation results and compare them with observational data, focusing on how well purely geometrical effects can explain observed $A_{V_{eff}}-R_{V_{eff}}$ relations. In Section 4, we analyze the link between dust properties and stellar mass, an environmental property related to SN Ia calibration, and to what extent it can be explained by our simulations. We also discuss how dust attenuation data could be used to quantitatively recover true extinction $R_{V}$ values, as well as the impacts of the source spectrum on the recovered dust properties. Finally, in Section 5, we summarize our main conclusions.
@@ 2 Methods
In this work, we make use of SKIRT, a powerful radiative transfer code that uses the Monte Carlo method (Camps & Baes, 2015, 2020) [^1] https://skirt.ugent.be/root/_home.html. This code relies on the simulation of the path of a number of photon packets emitted from a light source through a gridded astrophysical dusty medium. By taking into account both absorption and scattering effects, the code is able to stochastically simulate a view of the model from any orientation, as observed by a particular instrument. While it is also possible to simulate the contributions of dust emission, we do not do so, as it is not particularly relevant for the wavelengths in question in this work.
@@
The simulated astrophysical structure in a SKIRT model is made up of two systems, a light source and a dust medium system. The spectral energy distribution (SED) for the source system is defined based on the simple stellar populations (SSPs) described by Bruzual & Charlot (2003), assuming a Chabrier (2003) initial mass function (IMF), with a metallicity $Z=0.02$ and an age of $t_{age}=5$ Gyr. [^2] Given that SKIRT simulations output both the simulated observed flux and the total emitted flux, we can directly compute $A_{V_{eff}}$ and $R_{V_{eff}}$, meaning that the recovered dust properties are free from any degeneracies and that they are not impacted by the SED chosen for the source system.. To simulate aperture photometry observations, $10^{8}$ photon packets are used, while to simulate full-galaxy field observations (with a pixel size corresponding to a $25$ pc resolution), $10^{11}$ photon packets are used instead [^3] The value of $10^{11}$ photon packets, despite being quite high, was chosen because, for any lower number, severe stochastic fluctuations in $R_{V_{eff}}$ were present for the majority of the pixels.. We make use of SKIRT’s “ExtinctionOnly” mode given that, as stated above, we are not interested in dust emission.
By default, a Zubko et al. (2004) dust mix model is used for the dust medium, composed of bare silicate, graphite and polycyclic aromatic hydrocarbon (PAH) dust grains, with half of the PAH grains being neutral and half being ionized. A Mathis-Rumpl-Nordsiek (MRN) dust mix, composed of bare silicate and graphite (Mathis et al., 1977), is also used for a specific example, discussed in Section 3.2.2. The simulation outputs both the emitted and the observed flux, meaning that we can obtain accurate values for both $A_{V}$ and $R_{V}$ ($A_{V_{eff}}$ and $R_{V_{eff}}$, if we are in an attenuation scenario). While these two parameters are not ideal for a complete analysis of extinction and attenuation curves, given that they are restricted to the behavior of only two wavelengths, they nevertheless remain one of the most common parametrizations for dust effects, both in SN and galaxies.
We simulate observations from a variety of angles and apertures. We simulate global aperture photometry for an array of aperture SED detectors located at a distance of $10$ Mpc from the center of the model. The detectors are positioned at different angular positions, ranging from a face-on view of the galaxy at $0^{\circ}$ (along the z-axis) to an edge-on view at $90^{\circ}$ (along the xy-plane). The detector aperture radius is varied to examine different observing conditions. We also simulate full-galaxy field data for a detector that records the SED for each pixel in the field of view, also located at a distance of $10$ Mpc from the center of the model. A physical pixel size corresponding to a $25$ pc resolution is used.
2.1 Point Source/Dust Cloud Model
@@
Based on Baes et al. (2000), the dust cloud system density profile is described by a flattened spheroidal geometry given by Eq. 1:
@@
where $c=4000$ pc is the scale length. The normalization $\rho_{0}$ is varied to produce different optical depths in the dust system.

### v2

# Assessing differences between local dust attenuation and point source extinction within the same galactic environments

@@ Abstract
Methods. We employed SKIRT, a radiative transfer code, to simulate observations of point sources embedded in dust clouds, as well as spiral and elliptical galaxies. We examined various galaxy morphologies, inclinations, and instrument apertures.
Results. We find that in galaxies the scattering of light into the line of sight and the presence of sources at different depths within the galaxy make attenuation fundamentally different from extinction. For a medium with an intrinsic extinction curve slope of $R_{V}=3.068$, we recover effective attenuation curve slopes, $R_{V_{eff}}$, ranging from $0.5$ to $7$, showing that the two are not analogous, even for local resolved observations. We find that $R_{V_{eff}}$ greatly depends on dust density, galaxy morphology, and inclination, the latter being the most significant. A single simulated galaxy, viewed from different angles, can qualitatively reproduce the well-known relation between attenuation strength, $A_{V_{eff}}$, and $R_{V_{eff}}$ observed for star-forming galaxies. An increase in dust density leads to higher $R_{V_{eff}}$ across all inclinations, which, assuming a correlation between stellar mass and dust density, explains the increase in $R_{V_{eff}}$ with mass observed in star-forming galaxies. We cannot explain the observed differences in $R_{V_{eff}}$ between star-forming (higher $R_{V_{eff}}$) and quiescent (lower $R_{V_{eff}}$) high-mass galaxies, suggesting intrinsically lower $R_{V}$ values for ellipticals.
Conclusions. We conclude that highly attenuated regions of simulated face-on galaxies yield $R_{V_{eff}}$ within a $10\%$ error of the intrinsic extinction $R_{V}$ of the medium, allowing different dust types to be distinguished. For edge-on spirals, the median $R_{V_{eff}}$ for low $A_{V_{eff}}$ regions appears to better approximate the extinction $R_{V}$.
@@ 1 Introduction
Type Ia supernovae (SNe Ia) have long served as fundamental distance indicators for observational cosmology. They have contributed to the measurement of the accelerated expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999) and to the constraint of various cosmological parameters, such as the Hubble constant, $H_{0}$, (e.g., Dhawan et al., 2018; Riess et al., 2022; Galbany et al., 2023). However, when comparing the values of $H_{0}$ obtained via SNe Ia and other distance indicators of the distance ladder with those obtained from early-time probes, a tension of $\sim 5\sigma$ has been reported (Planck Collaboration et al., 2020; Riess et al., 2020). Although several cosmological model extensions have been proposed as the solution to this tension (e.g., Karwal & Kamionkowski, 2016; Zhao et al., 2017; Raveri, 2020), the possibility that some systematic error is at least partially responsible has not yet been fully discarded. In fact, recent results show that by using an inverse distance ladder whereby SN distances are calibrated via baryon acoustic oscillation (BAO) data, an $H_{0}$ compatible with early-time measurements can be recovered (Camilleri et al., 2025), pointing to the fact that, from the SN Ia point of view, the tension might originate in the standard local distance ladder (Wojtak & Hjorth, 2024).
@@
The process of SN Ia standardization for cosmology relies on a number of empirical corrections based on observed color-luminosity and light-curve shape-luminosity (Phillips, 1993; Tripp, 1998) relations, which have usually been assumed to be universal, i.e., the same for all SNe Ia (Guy et al., 2010). However, SNe Ia originating in high-mass galaxies have been shown to be more luminous than those originating in low-mass galaxies after the shape and color corrections (Sullivan et al., 2010; Kelly et al., 2010; Lampeitl et al., 2010). This effect is known as the “mass step” and it could hint at the existence of a systematic bias in the standardization, most likely related to the incomplete treatment of the previously mentioned color-luminosity relation, which could potentially influence SN Ia cosmological results.
@@
On top of its assumption of validity for all SNe Ia, the color-luminosity relation also combines two different physical effects in the same linear relation: i) intrinsic color-luminosity relations, i.e., fainter SNe Ia are redder; and ii) extrinsic dust effects, i.e., more dust extinction leads to fainter magnitudes and redder SNe through selective absorption. González-Gaitán et al. (2021) propose the existence of two separate populations of SNe Ia as an explanation for the mass step, with each having different intrinsic color-luminosity properties and dust environment characteristics. Brout & Scolnic (2021) suggest that host galaxy dust differences are solely responsible for the effect and define two separate $R_{V}$ populations for low- and high-mass galaxies, finding significantly different values for each one and an overall reduction in the mass step and Johansson et al. (2021) recover no mass step when using an individual best-fit $R_{V}$ for each standardized SN Ia. Similarly, Wiseman et al. (2022) find that using separate $R_{V}$ populations for younger and older galaxies results in a similar improvement, while Popovic et al. (2024) conclude that the mass step can, in part, be attributed to luminosity differences related to the age of the SN Ia progenitor and to its local environment. Duarte et al. (2023) find that SN host galaxy properties can be used to define a “dust step,” which nevertheless fails to fully account for the mass step. It has also been shown that apparent galactic dust properties can vary considerably between different galaxies and that they strongly correlate with environmental properties such as the stellar mass and the stellar age of the population in question (e.g., Garn & Best, 2010; Zahid et al., 2013; Salim et al., 2018; Duarte et al., 2023). As such, assuming a standard universal dust effect for SN Ia might prove inadequate in correcting the color-luminosity relation and could ultimately be at least partially responsible for the observed mass step.
@@
In an effort to improve SN Ia standardization, many methodologies have recently been proposed (e.g., Brout & Scolnic, 2021; Wiseman et al., 2022; Duarte et al., 2023), relying on the assumption that galactic dust properties can be used as an adequate proxy for SN Ia extinction. These approaches, while in some cases presenting a degree of success in reducing the Hubble residuals for the calibrated SN, might ultimately prove to be incomplete, given that attenuation and extinction are not the same. While they appear similar, the reddening laws describing the effects of dust on SN and other point source observations, also known as extinction laws, are fundamentally different from the reddening laws affecting integrated light observations such as galaxies, also known as attenuation laws (Krügel, 2009; Salim & Narayanan, 2020). This also means that the parameters usually used to describe dust effects, while sharing the same nomenclature in most of the literature, must be interpreted with care and in different ways.
@@
As light travels through the interstellar medium (ISM) and interacts with dust particles, it experiences absorption and scattering effects, which are especially prevalent for blue wavelengths (Calzetti et al., 1994). These effects lead both to a dimming and a reddening of the original spectral energy distribution, which can be parameterized by an empirical reddening law (e.g., Cardelli et al., 1989; Fitzpatrick, 1999; Calzetti, 2001).
@@
For a point source, the total extinction for the $V$ band, $A_{V}$, which serves as a normalization for the extinction curve, describes the dust column density in the optical path of the observed light. Likewise, the color excess for the $B-V$ color index, $E(B-V)$, refers only to the reddening experienced along that single line of sight. Therefore, the total-to-selective extinction ratio, $R_{V}=A_{V}/E(B-V)$, which describes the slope of the extinction curve, is only a function of the optical properties of the dust grains themselves (Draine, 2011).
@@
For an extended stellar population or a galaxy, this picture is complicated as a result of two primary factors. On the one hand, the necessity of using a larger aperture for galaxy observations makes the light scattered back into the line of sight a relevant component of the overall observed flux, with the strength of this scattered component being dependent on the morphology and inclination of the galaxy, as was demonstrated by Chevallard et al. (2013). Those authors show that galaxy inclination can drastically impact attenuation properties, meaning that identical galaxies viewed from different angles can produce wildly different attenuation curve strengths and slopes. On the other hand, the interspersion of stars and dust lanes leads to varying column densities inside the galaxy, along with different stellar light contributions. Viaene et al. (2017) found that the thicker the dust lane, the more the emission is dominated by stars located in the outermost regions of the galaxy, which are not meaningfully attenuated.
@@
Although this has long been known, it has only recently permeated the SN field. Chevallard et al. (2013) and Viaene et al. (2017) show that a deeper knowledge of the underlying galaxy geometry is needed to properly describe dust attenuation. Duarte et al. (2023) show that the SN Ia standardization returns much higher Hubble residuals when galactic dust properties are used to directly constrain the color luminosity correction. Popovic et al. (2024) conclude that dust in a SN line of sight cannot be accurately described using its host galaxy dust properties.
In this work we present a comprehensive analysis of how point source extinction relates to galactic attenuation. We rely on radiative transfer simulations to systematically study how factors such as instrument aperture and galaxy inclination, morphology and dust composition influence the recovered values of $A_{V_{eff}}$ and $R_{V_{eff}}$ and how these relate to the real optical properties of the medium. More specifically, we seek to understand under which conditions, if any, a host galaxy dust measurement can be said to truly reflect the extinction experienced by a local point source, such as a SN, and therefore be used as its proxy. We use the radiative transfer code SKIRT (Camps & Baes, 2015, 2020) to simulate observations for both point sources and a variety of galaxies (spiral and elliptical), observed from multiple angles and with differing apertures, the details of which are presented in Section 2. In Section 3, we present and discuss our simulation results and compare them with observational data, focusing on how well purely geometrical effects can explain observed $A_{V_{eff}}-R_{V_{eff}}$ relations. In Section 4, we analyze the link between dust properties and stellar mass, an environmental property related to SN Ia standardization, and to what extent it can be explained by our simulations. We also discuss how dust attenuation data could be used to quantitatively recover true extinction $R_{V}$ values, as well as the impacts of the source spectrum on the recovered dust properties. Finally, in Section 5, we summarize our main conclusions.
@@ 2 Methods
In this work, we made use of SKIRT, a powerful radiative transfer code that uses the Monte Carlo method (Camps & Baes, 2015, 2020) [^1] https://skirt.ugent.be/root/_home.html. This code relies on the simulation of the path of a number of photon packets emitted from a light source through a gridded astrophysical dusty medium. By taking into account both absorption and scattering effects, the code is able to stochastically simulate a view of the model from any orientation, as observed by a particular instrument. While it is also possible to simulate the contributions of dust emission, we did not do so, as it is not particularly relevant for the wavelengths in question in this work.
@@
The simulated astrophysical structure in a SKIRT model is made up of two systems, a light source and a dust medium system. The spectral energy distribution (SED) for the source system was defined based on the simple stellar populations (SSPs) described by Bruzual & Charlot (2003), assuming a Chabrier (2003) initial mass function (IMF), with a metallicity $Z=0.02$ and an age of $t_{age}=5$ Gyr. [^2] Given that SKIRT simulations output both the simulated observed flux and the total emitted flux, we can directly compute $A_{V_{eff}}$ and $R_{V_{eff}}$, meaning that the recovered dust properties are free from any degeneracies and that they are not impacted by the SED chosen for the source system.. To simulate aperture photometry observations, $10^{8}$ photon packets were used, while to simulate full-galaxy field observations (with a pixel size corresponding to a $25$ pc resolution), $10^{11}$ photon packets were used instead [^3] The value of $10^{11}$ photon packets, despite being quite high, was chosen because, for any lower number, severe stochastic fluctuations in $R_{V_{eff}}$ were present for the majority of the pixels.. We made use of SKIRT’s “ExtinctionOnly” mode given that, as was stated above, we are not interested in dust emission.
By default, a Zubko et al. (2004) dust mix model was used for the dust medium, composed of bare silicate, graphite, and polycyclic aromatic hydrocarbon (PAH) dust grains, with half of the PAH grains being neutral and half being ionized. A Mathis-Rumpl-Nordsiek (MRN) dust mix, composed of bare silicate and graphite (Mathis et al., 1977), was also used for a specific example, discussed in Section 3.2.2. The simulation outputs both the emitted and the observed flux, meaning that we can obtain accurate values for both $A_{V}$ and $R_{V}$ ($A_{V_{eff}}$ and $R_{V_{eff}}$, if we are in an attenuation scenario). While these two parameters are not ideal for a complete analysis of extinction and attenuation curves, given that they are restricted to the behavior of only two wavelengths, they nevertheless remain one of the most common parametrizations for dust effects, both in SN and galaxies.
We simulated observations from a variety of angles and apertures. We simulated global aperture photometry for an array of aperture SED detectors located at a distance of $10$ Mpc from the center of the model. The detectors were positioned at different angular positions, ranging from a face-on view of the galaxy at $0^{\circ}$ (along the z axis) to an edge-on view at $90^{\circ}$ (along the R plane). The detector aperture radius was varied to examine different observing conditions. We also simulated full-galaxy field data for a detector that records the SED for each pixel in the field of view, also located at a distance of $10$ Mpc from the center of the model. A physical pixel size corresponding to a $25$ pc resolution was used.
2.1 Point source in a dust cloud model
@@
Based on Baes et al. (2000), the dust cloud system density profile was described by a flattened spheroidal geometry given by Eq. 1:
@@
where $c=4000$ pc is the scale length. The normalization, $\rho_{0}$, was varied to produce different optical depths in the dust system.